A motorboat traveling with the current can go 24 m. in 2 hrs. Against the current it takes 3 hrs. to go with the same distance. Find the rate of the motorboat in calm water and the rate of the current?
Answer by mananth(16949)

against current x- 3.00 hours
with current x+ 2.00 hours

Distance = same= 24
Boat speed =x
current speed =y
t=d/r
24/(x-y)=3.00
3(x-y)= 24.00
3x-3y=24
x-y=8 ....................1
24/(x+y)=2.00
2.00(x+y)= 24
x+y=12 ...............2
add 1 & 2
2x=20
x=10 boat speed in still water
Plug value of x in 1
x-y=8
10-y=8
y=2 mph current speed